harliii:

An automobile engineer is redesigning a conical chamber that was originally specified to be 12 inches long with a circular base of diameter 5.7 inches. In the new design, the chamber is scaled by a factor of 1.5.

harliii:

The volume of the original chamber is 102.02 248.04 306.06 412.08 cubic inches. This is 242.47 344.49 432.16 448.21 cubic inches less than the volume of the new chamber

harliii:

those are the choices

Shadow:

Do you know the formula for the volume of a cone?

Shadow:

\[V_{cone} = \pi r^2 \frac{ h }{ 3 }\] Where r = radius and h = height

Shadow:

Let me know how you would approach this problem

harliii:

ik that

harliii:

idk how to do it

Shadow:

Solve for the volume of the first cone, then multiply it by the factor of 1.5 to get the new cone.

harliii:

i got the first one now what is the second one

harliii:

@shadow

Shadow:

More accurately, multiply the measurements by 1.5 to get the new cone

harliii:

i didn't get any of the answer choice

Shadow:

Let me work it out

harliii:

kk

harliii:

did u get the answer

Shadow:

\[V_{cone} = \pi r^2 \frac{ h }{ 3 }\] r = d/2 r = 5.7/2 r = 2.85 \[V_{1} = \pi r^{2}_{1} \frac{ h_{1} }{ 3 }\] r1 = 2.85 h1 = 12 \[V_{1} = \pi (2.85^{2})( \frac{ 12 }{ 3 })\] \[V_{1} = \pi(8.1225)(4) = 102.1\] Close enough to 102.02, they probably rounded something differently. \[V_{2} = \pi r^{2}_{2} \frac{ h_{2} }{ 3 })\] \[r_{2} = r_{1} \times 1.5\] \[r_{2} = 2.85 \times 1.5 = 4.275\] \[h_{2} = h_{1} \times 1.5 \] \[h_{2} = 12 \times 1.5 = 18\] \[V_{2} = \pi (4.275)^2 (\frac{ 18}{ 3 })\] \[V_{2} = \pi (18.275625)(6) = 344.49\]

Shadow:

Legend: V1 = volume of cone 1 r1 = radius of cone 1 h1 = height of cone 1 I am sure you get the idea. Same concept applies for the second cone. I did this so you know how to properly account for increasing the cone by a scale factor of 1.5.

Shadow:

Moved to the Mathematics section.